draw the graph of: y=f(x)=x^2+8x+7
Answer by ewatrrr(24785)

 
Hi,
f(x)=x^2+8x+7 = (x+4)^2 - 16 + 7 = (x+4)^2 -9
f(x)= 1(x+4)^2 -9 Parabola opening Upward 1 > 0, V(-4,-9)
Lineof Symmetry x = -4, P(0,7)& (-8,7) on Parabola